# Seeing the Outer Edge of the Infant Type Ia Supernova 2024epr in the Optical and Near Infrared

### v1

@@ Abstract
We present optical-to-near infrared (NIR) photometry and spectroscopy of the Type Ia supernova (SN Ia) 2024epr, including NIR spectra observed within two days of first light. The early-time optical spectra show strong, high-velocity Ca and Si features near rarely-observed velocities at 0.1, and the NIR spectra show a C i “knee.” Despite early-time, high-velocity features, SN 2024epr evolves into a normal SN Ia, albeit with stronger peak-light Ca absorption than other SNe Ia with the same light curve shape. Although we infer a normal decline rate, mag, from the light-curve rise, SN 2024epr is a Branch “cool” object and has red early-time colors ( mag at days). The high velocities point to a density enhancement in the outer layers of the explosion, predicted by some models, but thick-shell He-detonation models do not match the smoothly rising light curve or lack of He in our early-time NIR spectra. No current models (e.g., delayed detonation or thin He shell double detonation) appear to reproduce all observed properties, particularly the unusual early-time colors. Such constraints are only possible for SN 2024epr from the earliest optical and NIR observations, highlighting their importance for constraining SN Ia models. Finally, we find several other SNe Ia with intermediate mass elements at 30 000 km s-1 within days after the explosion that evolve into otherwise normal SNe Ia at peak light, suggesting the early-time spectra of SNe Ia may hide a broad diversity of observational characteristics.
@@
Type Ia supernovae (SNe Ia) are oft-occurring (Desai et al., 2024) astronomical transients associated with the thermonuclear explosions of carbon-oxygen white dwarf stars (CO WDs; Hoyle & Fowler 1960). They produce a significant fraction of the intermediate mass and Fe-group elements in the Universe (Raiteri et al., 1996; Matteucci & Recchi, 2001) and enable precise distance determinations (Phillips, 1993) useful for constraining cosmological parameters (Riess et al., 1998; Perlmutter et al., 1999; Burns et al., 2018; Jones et al., 2022; Riess et al., 2022). Despite the many SN Ia data sets (e.g., Hicken et al., 2009; Ganeshalingam et al., 2010; Brown et al., 2014; Krisciunas et al., 2017; Holoien et al., 2017a, b, c; Foley et al., 2018; Holoien et al., 2019; Jones et al., 2019; Phillips et al., 2019; Tucker et al., 2020; Fausnaugh et al., 2021; Jones et al., 2021; Scolnic et al., 2022; Fausnaugh et al., 2023; Neumann et al., 2023; Peterson et al., 2023; Morrell et al., 2024; Do et al., 2025) and theoretical models (e.g., Nomoto et al. 1984; Thielemann et al. 1986; Khokhlov 1991; Woosley & Weaver 1994; Hoeflich et al. 1995; Hoeflich & Khokhlov 1996; Iwamoto et al. 1999; Röpke & Niemeyer 2007; Shen et al. 2010; Woosley & Kasen 2011; Pakmor et al. 2012; Seitenzahl et al. 2013; Fink et al. 2014), the progenitor systems and explosion mechanisms of SNe Ia are not yet fully understood, and fundamental, qualitative questions remain unanswered (reviews include Maoz et al. 2014; Jha et al. 2019; Liu et al. 2023).
While there is broad consensus and strong evidence (Hoyle & Fowler, 1960; Nugent et al., 2011) that SNe Ia originate from exploding CO WDs, many disparate models exist. Each model includes a progenitor scenario (i.e., the constituent objects before any explosion occurs) and an explosion mechanism (i.e., how the CO WD ignites thermonuclear runaway). The two leading progenitor scenarios are the single-degenerate (SD) scenario with a CO WD and a non-degenerate companion such as a main sequence or red giant star (e.g., Whelan & Iben, 1973; Nomoto, 1982b; Nomoto et al., 1997) and the double-degenerate (DD) scenario with at least two CO WDs or a CO WD and a He WD (e.g., Nomoto, 1980; Iben & Tutukov, 1984; Webbink, 1984). Other scenarios exist as well, such as the core-degenerate (CD) scenario involving a CO WD and the degenerate CO core of an asymptotic giant branch star (AGB star; e.g., Hoeflich & Khokhlov, 1996), but are most often invoked only to explain highly peculiar SNe Ia (e.g., 2003fg-like SNe Ia Lu et al., 2021; Ashall et al., 2021).
@@
In the NIR, spectroscopic features have shallower optical depths than in the optical (Wheeler et al., 1998; Höflich et al., 2002) and thus probe material deeper in the ejecta than the optical regime at the same epoch. Unlike the optical, NIR features do not saturate (Ca ii H&K and NIR triplet and Si ii 6355 easily saturate, see, e.g., Hachinger et al., 2008), making them important diagnostics of SN Ia physics (see, e.g., Hsiao et al., 2019, and references therein). One example is tracing the distribution of unburnt carbon in the explosion, which differs between the deflagration (more unburnt carbon) and detonation (less unburnt carbon) models (e.g., Höflich et al., 2002; Kasen et al., 2009). Carbon is more easily measured with the less blended NIR C i features than optical C ii features, which suffer severe blending with Si ii (e.g., Hsiao et al., 2013, 2015; Marion et al., 2015; Wyatt et al., 2021). Additionally, there are two He features in the NIR (He i 1.083 and 2.058 m) that can be used to test predictions of He absorption made by some He-detonation models (e.g., Collins et al., 2023; Callan et al., 2024; Collins et al., 2024, although other models do not predict He features, e.g., Boos et al., 2021, 2024a). Unfortunately, few SNe Ia have optical and NIR spectral time series data within two days of the explosion (e.g., Hsiao et al., 2013, 2015; Pearson et al., 2024).
@@ 2 Data
The Zwicky Transient Facility (ZTF; Bellm et al., 2019) discovered SN 2024epr at on UT 2024-03-19 at 05:10:26 (MJD = 60388.2) with a -band magnitude of 19.4 mag (Sollerman, 2024a) and classified it as a pre-peak SN Ia (Sollerman, 2024b)111Note that Karambelkar et al. (2024) report an earlier spectrum within a day of discovery, but this spectrum was not uploaded as part of an official TNS classification report and remains proprietary.. The ZTF last non-detection was on UT 2024-03-10 at 03:26:47 (MJD = 60379.1) with a limiting -band magnitude of 18.9 mag. Figure 1 shows the location of SN 2024epr in NGC 1198 within the Pan-STARRS -band template image and a near-peak image taken by the Young Supernova Experiment (YSE; Jones et al., 2021; Aleo et al., 2023) as part of our photometric follow-up (see §2.2).
@@ 2.1 Host Galaxy
SN 2024epr exploded in NGC 1198, an elliptical (S0-̂, de Vaucouleurs et al., 1991) galaxy with a redshift (Huchra et al., 1999). The Milky Way line-of-sight extinction toward NGC 1198 is mag (Schlafly & Finkbeiner, 2011). To derive an approximate distance to NGC 1198, we use our assumed cosmological parameters, correct the redshift from the heliocentric to the cosmic microwave background frame, and apply a peculiar velocity correction using a density field derived from the 2M redshift catalog (Lavaux & Hudson, 2011). Our derived distance is Mpc ( mag), including a velocity error of 250 km s-1 in the error budget.
We use the Blast (Jones et al., 2024) tool222https://blast.scimma.org/. to acquire archival host-galaxy photometry and estimate the properties of NGC 1198. Blast measures galaxy photometry from the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006), the Wide-field Infrared Survey Explorer (Wright et al., 2010), the Sloan Digital Sky Survey (Fukugita et al., 1996; York et al., 2000), and the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al., 2016) using PSF-matched elliptical apertures and runs Prospector SED fits on the data (Leja et al., 2019; Johnson et al., 2021) with speed engagements from a simulation-based inference (SBI++) from Wang et al. (2023). Blast uses the Prospector- model with a non-parameteric star-formation history and parameters/priors (Jones et al., 2024, for more information, see Appendix A in).
@@ 2.2 Photometric Data
ATLAS observes the entire sky multiple times per night with four telescopes in Chile, South Africa, and Hawaiʻi (Tonry et al., 2018). It uses “cyan” and “orange” broadband filters that are approximately equivalent to + and + bandpasses, respectively. Most observations of SN 2024epr were taken with the orange () filter, which has a wavelength coverage of 560-820 nm (Tonry et al., 2018). Data were retrieved from the ATLAS Transient Science Server (Smith et al., 2020), and same-night data were stacked using a weighted average, excluding data affected by clouds.
@@
ZTF is a time-domain survey currently providing public, two-day cadence - and -bad data for the northern extragalactic sky. We obtained ZTF photometry using the ZTF forced photometry service (Masci et al., 2019)333https://irsa.ipac.caltech.edu/Missions/ztf.html..
Follow-up observerations from the Nickel and Lulin telescopes used CCD imagers. Point-spread-function photometry is measured from the images and calibrated by comparing field stars to the Pan-STARRS photometric catalog (Flewelling et al., 2020), using standard bias and sky-flat-field procedures.
NIR photometry was obtained with the Wide Field Camera (WFCAM; Casali et al. 2007; Hodgkin et al. 2009) mounted on the University of Hawai‘i owned and operated UKIRT 3.8m telescope. We again processed these data with Photpipe and calibrated them using 2MASS following Hodgkin et al. (2009) and Peterson et al. (2023). We used unforced, non-difference-imaged photometry from DOPHOT (Schechter et al., 1993); because SN 2024epr was bright and located near the outskirts of a smooth elliptical galaxy, a local background estimation and centroid determinations from individual images were considered sufficient for reliable photometry.
@@ 2.3 Spectroscopic Data
Seven NIR spectra were taken on the Gemini North 8m telescope using the Gemini Near-InfraRed Spectrograph (GNIRS; Elias et al. 2006b, a). Reductions were performed using standard PyPeit procedures (Prochaska et al., 2020a). An additional spectrum was taken using NASA’s InfraRed Telescope Facility (IRTF) with SpeX (Rayner et al., 2003) as part of program 2024A071 (PI: Hoogendam). Finally, the Keck Infrared Transients Survey (Tinyanont et al., 2024) obtained two post-peak transitional-phase NIR spectra with the Near-InfraRed Echellette Spectrometer on Keck II (NIRES; Wilson et al., 2004). The GNIRS, NIRES, and SpeX data were reduced with telluric corrections from an A0V star using the standard PypeIt (GNIRS and NIRES; Prochaska et al., 2020b) or Spextool (SpeX; (Cushing et al., 2004)) procedures. Table 3 catalogs our NIR spectroscopic observations.
@@ 3.1 Light Curves
Figure 2 shows the optical and NIR photometry for SN 2024epr, along with SuperNovae in object-oriented Python (SNooPy; Burns et al., 2011, 2014) fits (using the max_model, see following discussion) to the Pan-STARRS data. While we do not have -band data, SNooPy computes SED fits whose synthetic photometry is used to return parameters with respect to the band. Using EBV_model2, the SNooPy fit finds a host-galaxy extinction of mag and peak -band magnitude of 13.860.01 mag. However, the SNooPy EBV_model2 fit does not match the light curve of SN 2024epr well: it under-predicts the -band flux and declines slower than the -band light curve. The offset in the band may be due to the significant Ca absorption in the and bands, which could cause band-dependent shifts in color relative to a typical reddening law (see Section 4.1). We consider the high extinction measurement to be most likely due to the unusual colors of 2024epr or the fact that SNooPy may not fully model variation in SN intrinsic colors (Burns et al., 2014).
Therefore, we also ran a SNooPy fit with mag as a prior. We report the following results with (without) the host-galaxy extinction prior:
a -band magnitude peak on MJD (),
a color stretch of (), and
a distance modulus to NGC 1198 of () mag,
which corresponds to a distance of Mpc.
Finally, we perform fits using SNooPy’s max_model. This model returns results consistent with the EBV_model2 fits without the host-galaxy extinction prior. Since SN 2024epr has peculiar color evolution (see §3.2), we use the max_model values for , which is on MJD , and mag. The peak magnitudes are mag, mag, and mag, respectively.
@@
We emphasize that we do not directly measure . Rather, we infer it from fitting the rising light curve with SNooPy. To ensure the accuracy of our computed (and of the time of maximum light), we compare our phase-shifted light curves with SN 2020jgl (Galbany et al., 2025), another SN Ia with strong and high-velocity Ca ii absorption at early times (see §4 for spectral comparison). Figure 3 shows the overlaid light curves of SNe 2020jgl and 2024epr shifted to the rest-frame phase with respect to the epoch of peak -band magnitude and shifted by the difference in distance modulus between the two host galaxies. The light curves agree well using the SNooPy max_model peak times for SN 2024epr. SN 2020jgl has a similar value ( mag, Galbany et al., 2025) to our inferred value for SN 2024epr.
@@
We also fit the data with the SALT3 model (Kenworthy et al., 2021; Taylor et al., 2023) and see similar color discrepancies in both the early time and near-maximum data, with the -band flux in particular lower than the best-fit model. Again, we interpret this as an indication of stronger-than-normal Ca absorption affecting this band. The best-fit SALT parameters for SN 2024epr are a relatively red mag, a slightly fast (but uncertain) decline rate of , and a somewhat underluminous -band absolute magnitude that is approximately 0.7 mag fainter than the mean SN Ia at that redshift; however, adopting a peculiar velocity uncertainty of 250 km/s at the redshift of SN 2024epr gives an additional magnitude uncertainty 0.34 mag. Using the SALT nuisance parameters from Brout et al. (2022) gives a distance modulus to NGC 1198 of mag, closer than the SNooPy distance.
@@
Our inferred value is somewhat inconsistent with SN 2024epr being a Branch et al. (2006) ”cool” object (see 4.1) and the similarities in the NIR spectra to subluminous SNe Ia (see 5.3.3). In light of this, we also estimate based on the relationship between (the ratio of the Si ii 5972 and 6355 pEWs at peak light; see Nugent et al. 1995) and presented in Benetti et al. (2005), which gives mag777Benetti et al. (2005) do not fit the data, so we estimate by eye..
@@
Correcting for the distance and the Milky Way extinction,
we find that the -band absolute magnitude of SN 2024epr is mag (determined using the SNooPy SED fits). After making a standard correction for the SNooPy-derived EBV_model2 color, SN 2024epr is consistent with normal SNe Ia (Figure 4).

### v2

@@ Abstract
We present optical-to-near-infrared (NIR) photometry and spectroscopy of the Type Ia supernova (SN Ia) 2024epr, including NIR spectra observed within two days of first light. The early-time optical spectra show strong, high-velocity Ca and Si features near rarely-observed velocities at 0.1, and the NIR spectra show a C i “knee.” Despite early-time, high-velocity features, SN 2024epr evolves into a normal SN Ia, albeit with stronger peak-light Ca absorption than other SNe Ia with the same light curve shape. Although we infer a normal decline rate, mag, from the light-curve rise, SN 2024epr is a Branch “cool” object and has red early-time colors ( mag at days). The high velocities point to a density enhancement in the outer layers of the explosion, predicted by some models, but thick-shell He-detonation models do not match the smoothly rising light curve or apparent lack of He in our early-time NIR spectra. No current models (e.g., delayed detonation or thin He shell double detonation) appear to reproduce all observed properties, particularly the unusual early-time colors. Such constraints are only possible for SN 2024epr from the earliest optical and NIR observations, highlighting their importance for constraining SN Ia models. Finally, we identify several literature SNe Ia with intermediate mass elements at 30 000 km s-1 within days after the explosion that evolve into otherwise normal SNe Ia at peak light, suggesting the early-time spectra of SNe Ia may hide a broad diversity of observational characteristics.
@@
Type Ia supernovae (SNe Ia) are oft-occurring (e.g., Desai et al., 2024) astronomical transients associated with the thermonuclear explosions of carbon-oxygen white dwarf stars (CO WDs; Hoyle & Fowler 1960). They produce a significant fraction of the intermediate mass and Fe-group elements in the Universe (e.g., Raiteri et al., 1996; Matteucci & Recchi, 2001) and enable precise distance determinations (Phillips, 1993) useful for constraining cosmological parameters (Riess et al., 1998; Perlmutter et al., 1999; Burns et al., 2018; Jones et al., 2022; Riess et al., 2022). Despite the many SN Ia data sets (e.g., Hicken et al., 2009; Ganeshalingam et al., 2010; Brown et al., 2014; Krisciunas et al., 2017; Holoien et al., 2017a, b, c; Foley et al., 2018; Holoien et al., 2019; Jones et al., 2019; Phillips et al., 2019; Tucker et al., 2020; Fausnaugh et al., 2021; Jones et al., 2021; Scolnic et al., 2022; Fausnaugh et al., 2023; Neumann et al., 2023; Peterson et al., 2023; Morrell et al., 2024; Do et al., 2025) and theoretical models (e.g., Nomoto et al. 1984; Thielemann et al. 1986; Khokhlov 1991; Woosley & Weaver 1994; Hoeflich et al. 1995; Hoeflich & Khokhlov 1996; Iwamoto et al. 1999; Röpke & Niemeyer 2007; Shen et al. 2010; Woosley & Kasen 2011; Pakmor et al. 2012; Seitenzahl et al. 2013; Fink et al. 2014), the progenitor systems and explosion mechanisms of SNe Ia are not yet fully understood, and fundamental, qualitative questions remain unanswered (reviews include Maoz et al. 2014; Jha et al. 2019; Liu et al. 2023).
While there is broad consensus and strong evidence (Hoyle & Fowler, 1960; Nugent et al., 2011) that SNe Ia originate from exploding CO WDs, many disparate models exist. Each model includes a progenitor scenario (i.e., the constituent objects before any explosion occurs) and an explosion mechanism (i.e., how the CO WD ignites thermonuclear runaway). Two oft-invoked progenitor scenarios are the single-degenerate (SD) scenario with a CO WD and a non-degenerate companion such as a main sequence or red giant star (e.g., Whelan & Iben, 1973; Nomoto, 1982b; Nomoto et al., 1997) and the double-degenerate (DD) scenario with at least two CO WDs or a CO WD and a He WD (e.g., Nomoto, 1980; Iben & Tutukov, 1984; Webbink, 1984). Other scenarios exist as well, such as the core-degenerate (CD) scenario involving a CO WD and the degenerate CO core of an asymptotic giant branch star (AGB star; e.g., Hoeflich & Khokhlov, 1996), but are most often invoked only to explain highly peculiar SNe Ia (e.g., 2003fg-like SNe Ia Lu et al., 2021; Ashall et al., 2021).
@@
In the NIR, spectroscopic features have shallower optical depths than in the optical (Wheeler et al., 1998; Höflich et al., 2002) and thus probe material deeper in the ejecta than the optical regime at the same epoch. While optical features such as Ca ii H&K and NIR triplet and Si ii 6355 easily saturate (e.g., Hachinger et al., 2008)), NIR features do not, making them important diagnostics of SN Ia physics (see, e.g., Hsiao et al., 2019; Hoogendam et al., 2025; Müller-Bravo et al., 2025, and references therein). One example is tracing the distribution of unburnt carbon in the explosion, which differs between the deflagration (more unburnt carbon) and detonation (less unburnt carbon) models (e.g., Höflich et al., 2002; Kasen et al., 2009). Carbon is more easily measured with the less blended NIR C i features than optical C ii features, which suffer severe blending with Si ii (e.g., Hsiao et al., 2013, 2015; Marion et al., 2015; Wyatt et al., 2021). Additionally, there are two He features in the NIR (He i 1.083 and 2.058 m) that can be used to test predictions of He absorption made by some He-detonation models (e.g., Collins et al., 2023; Callan et al., 2024; Collins et al., 2024, although other models do not predict He features, e.g., Boos et al., 2021, 2024a). Unfortunately, few SNe Ia have optical and NIR spectral time series data within two days of the explosion (e.g., Hsiao et al., 2013, 2015; Pearson et al., 2024).
@@ 2 Data
The Zwicky Transient Facility (ZTF; Bellm et al., 2019) discovered SN 2024epr at on UT 2024-03-19 at 05:10:26 (MJD = 60388.2) with a -band magnitude of 19.4 mag (Sollerman, 2024a) and classified it as a pre-peak SN Ia (Sollerman, 2024b)111Note that Karambelkar et al. (2024) report an earlier spectrum within a day of discovery, but this spectrum was not uploaded as part of an official TNS classification report and remains proprietary.. The ZTF last non-detection was on UT 2024-03-10 at 03:26:47 (MJD = 60379.1) with a limiting -band magnitude of 18.9 mag. Figure 1 shows the location of SN 2024epr in NGC 1198 within the Pan-STARRS (Chambers et al., 2016) -band template image and a near-peak image taken by the Young Supernova Experiment (YSE; Jones et al., 2021) as part of our photometric follow-up (see §2.2).
@@ 2.1 Host Galaxy
SN 2024epr exploded in NGC 1198, an elliptical (S0-̂, de Vaucouleurs et al., 1991) galaxy with a redshift (Huchra et al., 1999). The Milky Way line-of-sight extinction toward NGC 1198 is mag (Schlafly & Finkbeiner, 2011). To derive an approximate distance to NGC 1198, we use our assumed cosmological parameters, correct the redshift from the heliocentric to the cosmic microwave background frame, and apply a peculiar velocity correction using a density field derived from the 2M redshift catalog (Lavaux & Hudson, 2011). Our derived distance is Mpc ( mag), including a velocity error of 250 km s-1 in the error budget. We adopt this distance.
We use the Blast (Jones et al., 2024) tool222https://blast.scimma.org/. to acquire archival host-galaxy photometry and estimate the properties of NGC 1198. Blast measures galaxy photometry from the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006), the Wide-field Infrared Survey Explorer (Wright et al., 2010), the Sloan Digital Sky Survey (Fukugita et al., 1996; York et al., 2000), and the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al., 2016) using PSF-matched elliptical apertures and runs Prospector SED fits on the data (Leja et al., 2019; Johnson et al., 2021) with speed engagements from a simulation-based inference (SBI++) from Wang et al. (2023). Blast uses the Prospector- model with a non-parameteric star-formation history and parameters/priors (for more information, see Appendix A in Jones et al., 2024).
@@ 2.2 Photometric Data
ATLAS observes the entire sky multiple times per night with four telescopes in Chile, South Africa, and Hawaiʻi (Tonry et al., 2018). It uses “cyan” and “orange” broadband filters that are approximately equivalent to + and + bandpasses, respectively. Most observations of SN 2024epr were taken with the orange () filter, which has a wavelength coverage of 560-820 nm (Tonry et al., 2018). Data were retrieved from the ATLAS Transient Science Server (Smith et al., 2020; Shingles et al., 2021), and same-night data were stacked using a weighted average, excluding data affected by clouds.
@@
ZTF is a time-domain survey currently providing public, two-day cadence - and -band data for the northern extragalactic sky. We obtained ZTF photometry using the ZTF forced photometry service (Masci et al., 2019)333https://irsa.ipac.caltech.edu/Missions/ztf.html..
Follow-up observations from the Nickel and Lulin telescopes used CCD imagers. Point-spread-function photometry is measured from the images and calibrated by comparing field stars to the Pan-STARRS photometric catalog (Flewelling et al., 2020), using standard bias and sky-flat-field procedures.
After uploading to arXiv, two of us (R.P. and P.C.) contacted the other authors with additional images taken with the Corrected Dall-Kirkham 24” and Ritchey-Chretien 32” telescopes by R.P. at the Post Observatory. The images were taken with Sloan filters produced by Astrodon with Pan-STARRS archival images as the template for image subtraction. The photometric calibration is against the SDSS magnitudes transformed from the Pan-STARRS photometric catalog (Flewelling et al., 2020).
NIR photometry was obtained with the Wide Field Camera (WFCAM; Casali et al. 2007; Hodgkin et al. 2009) mounted on the University of Hawai’i-owned and operated UKIRT 3.8m telescope. We again processed these data with Photpipe and calibrated them using 2MASS following Hodgkin et al. (2009) and Peterson et al. (2023). We used unforced, non-difference-imaged photometry from DOPHOT (Schechter et al., 1993). Because SN 2024epr was bright and located near the outskirts of a smooth elliptical galaxy, a local background estimation and centroid determinations from individual images were considered sufficient for reliable photometry.
@@ 2.3 Spectroscopic Data
Seven NIR spectra were taken on the Gemini North 8m telescope using the Gemini Near-InfraRed Spectrograph (GNIRS; Elias et al. 2006b, a). Reductions were performed using standard PyPeit procedures (Prochaska et al., 2020a). An additional spectrum was taken using NASA’s InfraRed Telescope Facility (IRTF) with SpeX (Rayner et al., 2003). Finally, the Keck Infrared Transients Survey (Tinyanont et al., 2024) obtained two post-peak transitional-phase NIR spectra with the Near-InfraRed Echellette Spectrometer on Keck II (NIRES; Wilson et al., 2004). The GNIRS, NIRES, and SpeX data were reduced with telluric corrections from an A0V star using the standard PypeIt (GNIRS and NIRES; Prochaska et al., 2020b) or Spextool (SpeX; (Cushing et al., 2004)) procedures. Table 3 catalogs our NIR spectroscopic observations.
@@ 3.1 Light Curves
Figure 2 shows the optical and NIR photometry for SN 2024epr, along with SuperNovae in object-oriented Python (SNooPy; Burns et al., 2011, 2014) fits (using the max_model, see following discussion) to the Post Observatory and Pan-STARRS data. While we do not have -band data near maximum light, SNooPy computes SED fits whose synthetic photometry is used to return parameters with respect to the band. Using EBV_model2 with mag as a prior (from the max_model fit described below), the SNooPy fit finds a host-galaxy extinction of mag and peak -band magnitude of 13.800.01 mag.
The SNooPy EBV_model2 model fits the light curve of SN 2024epr reasonably well in the bands, with worse fits in the bands. The issues in the bands may be due to the significant Si and Ca absorption, respectively, which could cause band-dependent shifts in color relative to a typical reddening law (see Section 4.1). We consider the high extinction measurement to be most likely due to the unusual colors of 2024epr or the fact that SNooPy may not fully model variation in SN intrinsic colors (Burns et al., 2014).
Since SN 2024epr has peculiar color evolution (see §3.2), we report the max_model values for
, which is on MJD ,
mag, and
The peak magnitudes are mag, mag, and mag, respectively.
Finally, the ATLAS last non-detection is notably deep. The 3 upper limit is 19.40 mag and the 5 upper limit is 18.85 mag. Correcting for the Milky Way extinction ( mag) and host-galaxy distance, this yields 3 and 5 absolute magnitude limits of mag and mag, respectively.
@@
We emphasize that we do not directly measure or because our data do not extend to the necessary time scales to do so (15 days and 20-40 days, respectively; Burns et al., 2014). Rather, we infer it from fitting the rising light curve with SNooPy. To ensure the accuracy of our computed (and of the time of maximum light), we compare our phase-shifted light curves with SN 2020jgl (Galbany et al., 2025), another SN Ia with strong and high-velocity Ca ii absorption at early times (see §4 for spectral comparison). Figure 3 shows the overlaid light curves of SNe 2020jgl and 2024epr shifted to the rest-frame phase with respect to the epoch of peak -band magnitude and shifted by the difference in distance modulus between the two host galaxies. The light curves agree well using the SNooPy max_model peak times for SN 2024epr. SN 2020jgl has a similar value ( mag, Galbany et al., 2025) to our inferred value for SN 2024epr.
@@
We also fit the data with the SALT3 model (Kenworthy et al., 2021; Taylor et al., 2023) and see similar color discrepancies in both the early-time and near-maximum data, with the -band flux in particular lower than the best-fit model. Again, we interpret this as an indication of stronger-than-normal Ca absorption affecting this band. The best-fit SALT parameters for SN 2024epr are a relatively red mag, a slightly fast (but uncertain) decline rate of , and a somewhat underluminous -band absolute magnitude that is approximately 0.7 mag fainter than the mean SN Ia at that redshift; however, adopting a peculiar velocity uncertainty of 250 km/s at the redshift of SN 2024epr gives an additional magnitude uncertainty 0.34 mag. Using the SALT nuisance parameters from Brout et al. (2022) gives a distance modulus to NGC 1198 of mag, closer than the SNooPy distance. For the rest of this work, we adopt the host-galaxy redshift distance from §2.1
@@
Our inferred value is somewhat inconsistent with SN 2024epr being a Branch et al. (2006) “cool” object (see 4.1) and the similarities in the NIR spectra to subluminous SNe Ia (see 5.3.3). In light of this, we also estimate based on the relationship between (the ratio of the Si ii 5972 and 6355 pEWs at peak light; see Nugent et al. 1995) and presented in Benetti et al. (2005), which gives mag777Benetti et al. (2005) do not fit the data, so we estimate by eye..
@@
Correcting for the distance and the Milky Way extinction, we find that the -band absolute magnitude of SN 2024epr is mag.
We correct for SN color by adopting the EBV_model2 host-galaxy extinction and , since the median for SNe Ia is known to be lower than that of the Milky Way (e.g., Burns et al., 2014). This is likely not from actual host-galaxy extinction but rather reflects the known empirical correlation between SN color and luminosity (e.g., Burns et al., 2014).
After this correction, SN 2024epr is consistent within uncertainties with normal SNe Ia (Figure 4). Its slightly subluminous location could be explained by a variety of factors, including a poor distance estimate to the host and the lack of a -band peak measured directly from the data, rather than inferred from an SED that may only superficially match that of SN 2024epr.